/

WLCG Tier1 Service Coordination Minutes - 1 March 2012

Attendance

Local:

Remote:

Action list review

Release update

Data Management & Other Tier1 Service Issues

Site Status Recent changes Planned changes
CERN CASTOR 2.1.11-9 (tapegateway active) + SRM-2.11 for all LHC instances; CASTOR 2.1.12-1 for PUBLIC instance This contains a new version of the xroot plug-in fixing potential problems (seen only on PUBLIC before the last upgrade: xroot-xcastor2fs_2112-1.1.0-1)
FTS: all nodes in SLC5 3.7.0-3
EOSATLAS and EOSCMS 0.1.2-2/xrootd-3.1; EOSALICE being set up
 

CASTOR 2.1.12 for all instances (ATLAS on Mar the 6th). All 3 remaining could be done the week after next.

ASGC CASTOR 2.1.11-6
SRM 2.11-0
DPM 1.8.2-3
23/2: All storage systems in downtime due to intervention on core switch None
BNL dCache 1.9.12.10 (Chimera, Postgres 9 w/ hot backup)
http (aria2c) and xrootd/Scalla on each pool
None None
CNAF StoRM 1.8.0 (Atlas, CMS, LHCb)    
FNAL dCache 1.9.5-23 (PNFS, postgres 8 with backup, distributed SRM) httpd=2.2.3
Scalla xrootd 2.9.7/3.1.0.osg
Oracle Lustre 1.8.6
EOS 0.1.1-12/xrootd 3.1.0.osg with Bestman 2.0.10
   
IN2P3 dCache 1.9.12-16 (Chimera) on core servers and pool nodes.
New hardware (more RAM, SSD disks) for Chimera and SRM servers (with SL6).
Postgres 9.1
Move from 1.9.5 to 1.9.12 (2nd golden release) nta
KIT dCache
atlassrm-fzk.gridka.de: 1.9.12-11 (Chimera)
cmssrm-fzk.gridka.de: head nodes 1.9.5-26 (Chimera), pool nodes 1.9.5-6 through -25
gridka-dcache.fzk.de: head nodes 1.9.5-26 (PNFS), pool nodes 1.9.5-24,-25
xrootd (version 20100510-1509_dbg)
   
NDGF dCache 2.1 (Chimera) on core servers. Mix of 1.9.13 and 2.0.1 on pool nodes.    
NL-T1 dCache 1.9.12-10 (Chimera) (SARA), DPM 1.7.3 (NIKHEF)    
PIC dCache 1.9.12-14 (last upgrade to patch release on 14-Dec); PNFS on Postgres 9.0    
RAL CASTOR 2.1.11-8
2.1.11-8 (tape servers)
SRM 2.11-1
Upgraded all CASTOR instances to 2.1.11-8 Migrate to new CASTOR DB hardware on 6 March
TRIUMF dCache 1.9.5-28 with Chimera namespace None None

Other site news

CASTOR news

CERN operations and development

EOS news

xrootd news

dCache news

StoRM news

FTS news

DPM news

LFC news

LFC deployment

Site Version OS, n-bit Backend Upgrade plans
ASGC NA NA NA NA
BNL 1.8.0-1 SL5, 64-bit Oracle None
CERN 1.8.2-0 64-bit SLC5 Oracle all servers are SLC5 64-bit virtual machines
CNAF 1.8.0-1 SL5 64-bit Oracle None
FNAL N/A     Not deployed at Fermilab
IN2P3 1.8.2-2 SL5 64-bit Oracle 11g  
KIT 1.8.2-2 SL5 64-bit Oracle 11g  
NDGF 1.7.4.7-1 Ubuntu 10.04 64-bit MySQL None
NL-T1 1.7.4-7 CentOS5 64-bit Oracle  
PIC 1.7.4-7 SL5 64-bit Oracle  
RAL 1.7.4-7 SL5 64-bit Oracle  
TRIUMF 1.7.3-1 SL5 64-bit MySQL None

Experiment issues

WLCG Baseline Versions

Status of open GGUS tickets

Review of recent / open SIRs and other open service issues

Conditions data access and related services

Database services

  • Experiment reports:
    • All databases upgraded to Oracle 11gR2 (11.2.0.3). Additional interventions to change the compatible parameter were necessary.
    • Latest Oracle security patches applied on all production databases.
    • Hw (storage) problems observed in RAC10. Several interventions already performed transparently. One issue this week due to high load also related to the storage problems. Being investigated by NetApp support.
    • CMS online database has been migrated today back to the CMS PIT.

  • Site reports:
Site Status, recent changes, incidents, ... Planned interventions
BNL Migration of Conditions database to 11gR2 (11.2.0.3) and a new hardware. In addition, latest CPU Jan 2012 patch was applied Decommissioning former Condition DB 10gR2 database Wed 29.02
CNAF    
KIT Old hw being decommissioned Maintenance downtime in the week 12.03-16.03: migration of Compass-LFC Schema to the new Hardware (FTS-DB) and upgrade to 11g. The FTS-DB will be down on 14.03 approximately between 9:00 14:00 CET.
IN2P3 Waiting for patch to fix issue with Oracle installer (SR 3-5225032577, bug 13731278)  
PIC FTS database was upgraded last tuesday, and at the same time, LFC of ATLAS database was completely stopped due to LFC migration to CERN. Currently all databases are in 11.2.0.3. There are no scheduled interventions.
RAL Applied the Jan PSU on the 3D system (11.2.0.3.1). Compatible parameter has not been changed accordingly since the intervention was announced as an at risk. Castor DB will be moved to new HW next week on Tuesday
SARA    
TRIUMF ntr  

  • Decommissioning of the 3D OEM.

AOB

-- AndreaSciaba - 29-Feb-2012

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2012-05-02 - AndreasPetzold
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback